[Lerftest] LERF Update

Kevin Jordan jordan at jlab.org
Tue Sep 15 07:36:13 EDT 2020

Started shift finishing up a few items left on MPS checkout, this took longer than expected & consumed the day shift. Swing shift had gun HVPS MPS interlock problem, this was fixed and then is was discovered that the Drive Laser Pulse Controller (DLPC) was not “talking” with the MPS (there are 8 fiber optic cables, when all are active CW is allowed, each one has a lower beam power limit). System expert was not able to come in to fix problem last night so it will be addressed first thing this AM.
Pulsed mode operations began to find remaining problems. There were also a couple viewer/camera issues that need repair. Beam threading to 1X will continue this morning.

Swing shift summary:
Gun HVPS controller repair in progress at start of shift, completed around 17:00. MPS checkout of gun completed.
DLPC SSG certification started. Certification punted until tomorrow morning when hardware was found to behave inconsistent with what was expected.
EES-Rf contacted to assist with tuning 4L02-1.
Quads reloaded from March 10 all-save.
Pulsed mode beam sent as far as 1F05, and all-save taken.
Shutdown script run and turned over to MCC for hot standby.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/lerftest/attachments/20200915/7464ac7e/attachment.html>

More information about the Lerftest mailing list